# CrossDehaze: Scaling Up Image Dehazing with Cross-Data Vision Alignment and Augmentation

### v1

@@ Abstract
In recent years, as computer vision tasks have increasingly relied on high-quality image inputs, the task of image dehazing has received significant attention. Previously, many methods based on priors and deep learning have been proposed to address the task of image dehazing. Ignoring the domain gap between different data, former de-hazing methods usually adopt multiple datasets for explicit training, which often makes the methods themselves be violated. To address this problem, we propose a novel method of internal and external data augmentation to improve the existing dehazing methodology. By using cross-data external augmentor. The dataset inherits samples from different domains that are firmly aligned, making the model learn more robust and generalizable features. By using the internal data augmentation method, the model can fully exploit local information within the images, thereby obtaining more image details. To demonstrate the effectiveness of our proposed method, we conduct training on both the Natural Image Dataset (NID) and the Remote Sensing Image Dataset (RSID). Experimental results show that our method clearly resolves the domain gap in different dehazing datasets and presents a new pipeline for joint training in the dehazing task. Our approach significantly outperforms other advanced methods in dehazing and produces dehazed images that are closest to real haze-free images. The code will be available at: https://github.com/wengzp1/ScaleUpDehazing
@@
Image Dehazing, Vision Alignment, Vision Augmentation, Remote Sensing, Self-Supervised Learning.
Figure 1: Our method compared to other dehazing methods in terms of flowchart and characteristics. From the perspective of performance and methodology, we apply Dehamer [1] and Trinity-Net [2] for comparison. Ignoring the domain gap between different data, former de-hazing methods usually adopt multiple datasets for explicit training, which often make the methods itself be violated. To address this issue, we call for the internal- and external knowledge can be further augmented with a self-supervised learning to fill up the domain gap.
@@ I Introduction
In the field of remote sensing images and natural images, haze usually leads to a decrease in image contrast and clarity. It makes the originally distinct boundaries of objects blurry, greatly affecting the visual effect of the image. Therefore, developing effective dehazing methods is of great significance for enhancing the application value of remote-sensing images. Previous dehazing methods mainly address hazy images by using atmospheric scattering models [3][4][5]. The model explains how light interacts with particulate matter in the atmosphere under hazy conditions, thereby affecting the imaging process of the images. The specific representation is as follows:
$I(x)=J(x)t(x)\oplus A(1-t(x)),$ (1)
@@
where x is a pixel position in the image, $J(x)$ represents the true haze-free image, and A is the global atmospheric light. $t(x)$ represents the medium transmission rate, which ranges from 0 to 1. Here, 0 indicates no transmission, meaning complete opacity, while 1 indicates complete transparency. $I(x)$ represents the hazy image obtained after the haze-free image undergoes atmospheric scattering.
[…]
Recently, many methods have emerged for use in the field of image dehazing. ROP [9] introduces a rank-one matrix to represent complex scenes as linear combinations of a series of simple basic elements and then uses these basic elements for image dehazing. Cycle-SNSPGAN [10] improves the quality of dehazed images by combining spectral normalization and patch discriminator techniques. DEFADE [11] achieves perceptual prediction of haze density by analyzing the statistical features of natural scenes and haze-sensitive features in the image. In addition, some deep learning methods train on a large number of hazy and haze-free image pairs, and use models such as CNNs or Transformers to learn the mapping from hazy images to clear ones. UHD [12] proposes a multi-guided bilateral learning method, which combines features from multiple scales and levels of images as well as bilateral filtering techniques to obtain dehazed images. Dehamer [1] introduces a transmission-aware 3D positional embedding module into the Transformer, providing relative positional information while incorporating priors related to haze density. Trinty-Net [2] consists of three parameter estimation networks: the haze thickness estimator network(S-Net), the ambient light estimator network(A-Net), and the medium transmission estimator network(T-Net). Obtaining these parameters further generates dehazed images.
@@
However, the aforementioned methods mostly consider the model perspective and rarely take the data aspect into account. The dehazing results obtained by the aforementioned methods still have limitations due to the narrow training data. Specifically, the training dataset lacks diversity in terms of scenes and types of haze, resulting in the model’s inability to achieve good robustness. Moreover, there is a domain gap between existing datasets, and directly using multiple datasets for training often results in limited performance improvement.
@@
To address the issue of data scarcity and uniqueness, we propose a novel method of external data augmentor and internal data augmentor to improve the dehazing model. Specifically, we first expand the dataset externally using a channel alignment method. Next, we apply strong-weak augmentation methods internally and perform self-supervised learning on the resulting images within the dataset. Finally, we compare our approach with the advanced dehazing works, and the performance metrics demonstrate the superiority of our method. As shown in Fig. 1, it can be observed that our method outperforms Dehamer and Trinity-Net in terms of quantitative metrics such as PSNR and SSIM from the obtained dehazed results. Additionally, the dehazing results produced by our method visually appear closer to real haze-free images. In summary, the contributions of this paper can be summarized as follows:
@@
    We proposed a cross-data external augmentor. The dataset inherits samples from different domains that are aligned by the external data augmenter, making the model learn more robust and generalizable features.
@@
    We proposed an internal data augmentor in a strong-weak fashion. The model can effectively explore local information within images. While the model focuses on the overall image externally, the local details of the image also be fully explored.
  • $\bullet$
    Experiments have shown that our method achieves state-of-the-art results. Our method outperforms other advanced methods in terms of PSNR and SSIM metrics.
@@ II RELATED WORK
II-A Remote-sensing Image Dehazing
In the fields of geography and remote sensing, image dehazing techniques play an important role in image preprocessing. Due to the poor visibility of remote sensing images or natural images acquired in hazy environments, it is visually easy to cause the loss of details [7, 13, 14, 15, 16]. Hazy images are not conducive to the related tasks of computer vision, thus causing a lot of challenges to a certain extent. For example, hazy images make it difficult for object detection models to accurately identify and localize objects in the images [17, 18]. In semantic segmentation tasks [19], the presence of haze makes scene boundaries that should be clear become blurred, leading to segmentation models struggling to differentiate between different objects or scene elements. In image recognition tasks [20], the presence of haze interferes with the model’s overall understanding of the image content.
The methods for dehazing remote sensing images can be mainly classified into two categories: those based on priors and those based on learning. Early, He et al. proposed the Dark Channel Prior (DCP) [7] method. They observed that in at least one color channel, there are always some pixels in local regions of the image with very low intensity. This prior is used to estimate the distribution of haze in the image. Color Attenuation Prior (CAP) [21] is a method based on physical models. It assumes that in a hazy environment, the color of a scene gradually attenuates as the depth increases. This prior can be used to estimate the depth map of the scene, thereby restoring the clear image. Color-lines [22] analyze the color information in the image to estimate the distribution of haze and depth information of objects in the scene. Haze-lines [23] involves a deep understanding and analysis of image features under hazy conditions. By identifying hazy regions in the image and decomposing and processing them, haze-lines aim to remove the effects of haze.
Recently, with the development of deep learning techniques, deep learning-based dehazing methods have gradually become a research hotspot. DehazeNet [24] utilizes the deep architecture of CNN to output the medium transmission map. Then, it uses the atmospheric scattering model to recover the haze-free image. MSCNN [25] combines a multi-scale information extraction subnetwork and a multi-scale residual dilated convolution module. It integrates features from different receptive fields to estimate the transmission map of hazy images. DCPDN [26] employs two generators, one for generating the transmission map and the other for estimating the atmospheric light. GFN [27] processes images through operations such as white balancing and utilizes an end-to-end trainable neural network with a structure similar to Unet for dehazing. Dehamer [1] introduces a three-dimensional positional embedding module into the Transformer, incorporating priors related to haze density to estimate the haze density in different spatial regions. Trinity-Net [2] consists of three parameter estimation networks, each estimating haze thickness, ambient light, and medium transmission, respectively.
II-B Data Augmentation
Data augmentation is an important technique [28] in deep learning, which mainly involves artificially increasing the amount of data by generating new data samples from existing data. Previous image data augmentation operations often increase the size of the dataset by altering the image shape or visual appearance. Common techniques include injecting noise, flipping, cropping, resizing, and color space operations [29, 30, 31]. Current data augmentation methods also include techniques based on image erasure. Cutout [32] is a data augmentation technique that involves randomly erasing subregions during training and filling them with 0 or 255 in the image. Random erasing [33] is an augmentation method that randomly erases subregions in the image, similar to cropping. However, it also randomly determines whether to mask and decides the aspect ratio and size of the masked region. The key idea of Hide-and-Seek [34] is to divide the image into randomly sized uniform squares and randomly delete a random number of squares. When important information is hidden, it forces the neural network to learn relevant features.
Recently, non-single-image mixed data augmentation methods have become increasingly common techniques in data augmentation. Mixup [35] combines two randomly selected images and their corresponding labels, where the mixing ratio is determined by a mixing coefficient $\alpha$. This method can improve the accuracy of the model. CutMix [36] addresses the problem of information and local feature loss. CutMix randomly selects a region from each image when mixing images. Then it combines them to form a mixed image after swapping these regions. Puzzle Mix [37] introduces a puzzle-based mixed data augmentation technique. This method focuses on flexibly utilizing key information and basic statistical data of images, aiming to break the neural network’s reliance on existing data augmentation. RandomMix [38] aims to enhance the generalization performance of the model. It randomly selects mixed augmentations from a set of augmentations and applies them to the image, allowing the model to encounter different samples. However, directly using mixed data augmentation methods is not suitable for image dehazing tasks. Using mixed data augmentation methods directly alters the visual content of the original image, potentially resulting in the loss of crucial information during the dehazing process.
In image dehazing tasks, data augmentation operations often need to be designed specifically for the characteristics of images under hazy conditions. To address this issue, we propose an external data augmentation and an internal data augmentation method. This method can enhance the model’s ability to handle hazy images and improve its robustness.
II-C Self-supervised Learning
Self-supervised learning(SSL) has achieved significant success in the field of computer vision [39, 40, 41, 42, 43, 44, 45]. It trains models by generating labels automatically, without the need for manually annotated datasets. Many researchers have proposed different solutions for creating pseudo-labels. For example, flipping images, cropping or merging image patches, etc. The key idea of self-supervised learning is to utilize the inherent structure and relationships within the data itself to learn effective feature representations. Self-supervised learning can be trained on large amounts of unlabeled data, which is highly valuable in fields like remote sensing image dehazing and natural image dehazing. This is because acquiring large amounts of labeled data is often expensive and time-consuming.
@@ III Methodology
As shown in Fig. 2, our proposed method consists of two parts: external data augmentor and internal data augmentor. Firstly, we propose an external data augmentation method to reduce the contrast differences between different hazy datasets. External data augmentor makes the augmented dataset approximate the target dataset in the color space. The augmented samples obtained by our method can better match the distribution of the target domain. To this end, dehaze networks trained with consistent yet more quantity of samples have the potential to show superior results.
Building upon the augmentation of different domain datasets externally, we further explore additional locally effective information by augmenting the dataset internally. The model becomes more robust by coordinating external and internal augmentation. For internal data augmentation, we introduce a weak-to-strong coordinated data augmentation self-supervised learning method to improve the local details of images. The model can focus more on the details of the images by calculating the difference between weakly and strongly augmented images.
Figure 2: Illustration of the proposed integrated framework for External-Augmentor with Auxiliary Datasets (External-Aug. with AD.) and Internal-Augmentor with Self-supervised Learning (Internal-Aug. with SSL.). Our method first performs augmentation on the external dataset to enrich dataset diversity. The external augmentor makes the auxiliary dataset be aligned with the target dataset in the color channels. Then, we apply the self-supervised learning method to augment the data internally. Specifically, the internal augmentor emphasizes attention to fine-grained details in images. By simultaneously augmenting the data externally and internally, our model achieves higher performance and robustness.
III-A Cross-data External Augmentor for Alignment
As described in Sec. II, traditional data augmentation methods often rely on simple geometric and color transformations. However, this approach to processing images generated from the original dataset lacks diversity. Additionally, there are often significant differences between different dehaze datasets in terms of color and contrast. Suppose we directly use several datasets for joint training, the corresponding experiments often show a degradation.
In particular, we propose an external-augmentor to address the domain gap between different datasets. The proposed external-augmentor achieves efficient domain adaptation without an additional training phase. Specifically, we first calculate the gamma values $\gamma_{R,G,B}$ for the three channels to be transferred to the target domain. Then, we perform gamma calibration on the images to be transferred, as follows:
@@ III-A Cross-data External Augmentor for Alignment
where $I_{R,G,B}$ and $O_{R,G,B}$ represent the input and output pixel intensities ([0,255]), respectively, and $\gamma$ is the gamma factor. The subscripts R, G, B represent the color channels, and the values of different channels are unique. Our method adjusts the RSID dataset to make it more similar to the NID dataset from color space. As shown in Fig. 3, the RSID images better be aligned with the RGB distribution of the NID dataset after external augmentation. By using external augmentation, our method effectively expands the dataset’s quantity and introduces external datasets, enriching the diversity of the dataset. Some of the transformed data is shown in Fig. 4. The first column shows the original RSID images, the second column shows the new RSID images after data preprocessing, and the third column shows the target dataset NID. As shown in Fig. 4, the RSID images become closer to the NID dataset after the external augmentor. To illustrate our method intuitively, we demonstrate the algorithmic flow of external augmentation. As shown in Algorithm 1, we apply the RSID dataset as an example to present the implementation of the channel correction.
Figure 3: A demonstration of domain gap between different dahazing datasets. We perform gamma correction on the RSID dataset based on the RGB values of the NID dataset. It can be observed that before correction, there is a significant difference between the distributions of the datasets in the RGB channels. After external data augmentation, the distributions of RSID images are closely aligned with those of NID images in the RGB channels.
@@
Algorithm 1 External-Augmentor for Data Alignment
@@
     $O_{CM-R}$=$\left(\frac{I_{R}}{255}\right)^{\frac{1}{\gamma_{R}}}$
     $O_{CM-G}$=$\left(\frac{I_{G}}{255}\right)^{\frac{1}{\gamma_{G}}}$
     $O_{CM-B}$=$\left(\frac{I_{B}}{255}\right)^{\frac{1}{\gamma_{B}}}$
@@
Then we can get RSID hazy images after channel migration transformation
Algorithm 2 Internal-Augmentor with Self-supervised Learning
@@
III-B Feature Extractor
As shown in Fig. 2, the dehazing network is composed of a 5-layer U-Net structure [46]. The network includes two downsampling and upsampling steps. In the network, the input hazy image passes through Dehaze Block modules and feature fusion modules to finally obtain the output dehazed image $\widehat{y}$. As shown in part B of Fig. 2, in the Dehaze Block module, the input feature maps undergo normalization and padding layers. Then they are projected into Q , K, and V (query, key, value) using linear layers, used to calculate attention scores [47]. Additionally, the Dehaze Block module applies multi-head self-attention within a window to handle relationships between image patches. It allows the dehazing network to consider correlations between different positions while extracting features. Thus, the model can learn structural information in the image more effectively to improve the model performance.

### v2

# Scaling Up Single Image Dehazing Algorithm by Cross-Data Vision Alignment for Richer Representation Learning and Beyond

@@ Abstract
In recent years, deep neural networks tasks have increasingly relied on high-quality image inputs. With the development of high-resolution representation learning, the task of image dehazing has received significant attention. Previously, many methods collect diverse image data for large-scale training to boost the performance on a target scene. Ignoring the domain gap between different data, former de-hazing methods simply adopt multiple datasets for explicit large-scale training, which often makes the methods themselves be violated. To address this problem, we propose a novel method of cross-data vision alignment for richer representation learning to improve the existing dehazing methodology. Specifically, we call for the internal- and external knowledge should be further adapted with a self-supervised manner to fill up the domain gap. By using cross-data external alignment, the datasets inherit samples from different domains that are firmly aligned, making the model learn more robust and generalizable features. By using the internal augmentation method, the model can fully exploit local information within the images, and then obtaining more image details. To demonstrate the effectiveness of our proposed method, we conduct training on the Natural Image Dataset (NID). Experimental results show that our method clearly resolves the domain gap in different dehazing datasets and presents a new pipeline for large-scale training in the dehazing task. Our approach significantly outperforms other advanced methods in dehazing and produces dehazed images that are closest to real haze-free images. The code will be available at: https://github.com/wengzp1/ScaleUpDehazing
@@
Image Dehazing, Vision Alignment, Large-scale, Representation Learning, Self-Supervised Learning.
@@ I Introduction
In the field of natural images, haze usually leads to a decrease in image contrast and clarity. It makes the originally distinct boundaries of objects blurry, greatly affecting the visual effect of the image. Therefore, developing effective dehazing methods is of great significance for enhancing the application value of remote-sensing images. Previous dehazing methods mainly address hazy images by using atmospheric scattering models [1][2][3]. The model explains how light interacts with particulate matter in the atmosphere under hazy conditions, thereby affecting the imaging process of the images. The specific representation is as follows:
$I(x)=J(x)t(x)\oplus A(1-t(x))$ (1)
@@
where $x$ is a pixel position in the image, $J(x)$ represents the true haze-free image, and $A$ is the global atmospheric light. $t(x)$ represents the medium transmission rate, which ranges from 0 to 1. Here, 0 indicates no transmission, meaning complete opacity. $I(x)$ represents the hazy image obtained after the haze-free image undergoes atmospheric scattering.
Early dehazing methods utilize prior knowledge of the image to separately estimate the haze parameters $t(x)$ and $A$, and then further use the atmospheric scattering model to obtain the haze-free image $J(x)$. ALC [4] proposes an image dehazing method that combines atmospheric light white balance correction, local light filtering, and high-altitude perspective priors to restore clarity to hazy images. While these methods are intuitive, they struggle to effectively remove haze in non-uniformly hazy images.
Recently, many methods have emerged for use in the field of image dehazing. MSPD-Net [5] proposes a progressively optimized multi-stage dehazing network. It uses a physics-based feature enhancement block and selective kernel feature fusion to achieve image dehazing. MDFEN [6] proposes an enhanced network that fuses multiscale depth information. This network integrates multiscale features from depth maps and hazy images. It also incorporates a context attention mechanism to improve dehazing performance in scenes with deep depth of field and heavy fog. Oval-Net [7] proposes an end-to-end dehazing network that incorporates spatial and channel attention mechanisms. The network efficiently processes image data within an encoder-decoder structure, improving dehazing performance without relying on the atmospheric scattering model. TOENet [8] proposes an enhancement network designed to improve image quality in low-visibility environments. It uses a multilayer perceptron-based channel correlation extraction module and an encoder-decoder architecture to achieve image dehazing and enhancement in hazy and sandstorm conditions. Cycle-SNSPGAN [9] improves the quality of dehazed images by combining spectral normalization and patch discriminator techniques. In addition, deep learning methods train on a large number of hazy and haze-free image pairs, and use models such as Transformers to learn the mapping from hazy images to clear ones. Dehamer [10] introduces a transmission-aware 3D positional embedding module into the Transformer, providing relative positional information while incorporating priors related to haze density.
@@
However, the aforementioned methods mostly consider the model perspective and rarely take the data itself aspect into account. With the rapid development of Transformer [11, 12], the demand for data has grown at the same time. Ignoring the domain gap between different data, former de-hazing methods usually adopt multiple datasets to a large model (e.g., Transformers), which often achieves unsatisfactory results. As shown in Fig. 1, the Dehamer [11] applies the narrow training data to train a Transformer. Specifically, the training dataset lacks diversity in terms of scenes and types of haze, resulting in the model’s inability to achieve good robustness. Moreover, Trinity-Net [13] ignores the domain gap between existing datasets, and directly using multiple datasets for training, which often results in limited performance improvement.
Figure 1: Comparison with state-of-the-arts in terms of flowchart and characteristics. With the rapid development of Transformer [11, 12], the demand for data has grown at the same time. However, the simple combination of multiple datasets achieves unsatisfactory results. To address this issue, we call for the internal- and external knowledge should be further augmented by vision alignment and self-supervised learning to perform an effective large-scale training.
Then, we compare existing self-supervised dehazing methods with our method. SLAdehazing [14] employs a self-supervised adaptation strategy, leveraging contrastive learning directly on a large-scale synthetic dataset of hazy-clean image pairs. In contrast, our method enhances external data through cross-dataset alignment, significantly improving the model’s robustness. Instead of guiding dehazing with depth information, as in DP-dehazing [15], our method utilizes gamma correction and Weak-to-Strong Augmentation method to align the distributions of the training data. In comparison to SGDRL [16], which focuses on disentangling features for dehazing, our method integrates cross-data alignment with self-supervision, emphasizing the robustness gained from training on diverse augmented data.
@@
To address the data scarcity and uniqueness in large-scale training, we propose a novel method of external- and internal- knowledge alignment and augmentation, to improve the dehazing model. Specifically, we first expand the dataset externally using a knowledge alignment method. The datasets inherit samples from different domains that are firmly aligned, making the model learn more robust and generalizable features. Next, we apply weak-to-strong self-supervised learning methods to learn richer representations. Additionally, the dehazing results produced by our method visually appear closer to real haze-free images. To this end, the contributions of this paper can be summarized as follows:
@@
    We expand the dataset externally using a knowledge alignment method. The dataset inherits samples from different domains that are aligned by the external data augmenter, making the model learn more robust and generalizable features.
@@
    We propose a Weak-to-Strong Augmentation method with self-supervised manner to learn richer representations. Compared with previous approaches, the proposed method learns more feature representations with spatial-invariant strategies and self-supervised learning.
Figure 2: Illustration of the proposed integrated framework for External-Augmentor with Auxiliary Datasets (External-Aug. with AD.) and Internal-Augmentor with Self-supervised Learning (Internal-Aug. with SSL.). In our method, we perform external augmentation on the target dataset to increase its diversity. The external augmentor makes the auxiliary dataset be aligned with the target dataset in the color channels. And gamma correction mainly affects the brightness and contrast of the images. Then, we apply the self-supervised learning method to augment the data internally. Specifically, the internal augmentor emphasizes attention to fine-grained details in images. By simultaneously augmenting the data externally and internally, our model achieves higher performance and robustness.
@@ II RELATED WORK
II-A Image Dehazing
In the fields of natural image processing, image dehazing techniques play an important role. Hazy images are not conducive to the related tasks of vision, thus causing a lot of challenges to a certain extent. Recently, with the development of deep learning techniques, deep learning-based dehazing methods have gradually become a research hotspot. MSPD-Net [5] introduces a multistage progressive dehazing network. It achieves effective image dehazing using a physical model-based feature enhancement block and selective kernel feature fusion. MDFEN [6] proposes a multiscale depth fusion enhancement network. Oval-net [7] introduces an end-to-end dehazing network based on spatial and channel attention mechanisms. TOENet [8] introduces a two-in-one low-visibility enhancement network. FSE-DIF [17] introduces a fast single-image dehazing algorithm based on artificial multi-exposure image fusion. HECB [18] introduces a contrast enhancement algorithm based on the atmospheric scattering model. DehazeNet [19] utilizes the deep architecture of CNN to output the medium transmission map. Then, it uses the atmospheric scattering model to recover the haze-free image. Liu et al [20] first employ a gamma correction on non-homogeneous dehazing task, which achieves a impressive results and gradually become a basic operation in dehazing task. Dehamer [10] introduces an three-dimensional positional embedding module into the Transformer, incorporating priors related to haze density to estimate the haze density in different spatial regions. Previous methods ignore the data scarcity and uniqueness in large-scale training, which often results in limited performance improvement.
II-B Self-supervised Learning
Self-supervised learning(SSL) has achieved significant success in the field of computer vision [21, 22, 23, 24, 25, 26, 27, 28, 29]. It trains models by generating labels automatically, without the need for manually annotated datasets. The key idea of self-supervised learning is to utilize the inherent structure and relationships within the data itself to learn effective feature representations. Self-supervised learning can be trained on large amounts of cheap data, which is highly valuable in fields like image dehazing and natural image dehazing. Recently, many self-supervised learning-based methods have been applied to dehazing tasks. SLAdehazing [14] removes haze by pre-training on synthetic hazy-clean image pairs and adapting to real hazy images through contrastive learning. DP-dehazing [15] designs a depth-guided hybrid attention Transformer module, leveraging depth information to enhance dehazing performance. SGDRL [16] achieves unsupervised and semi-supervised single image dehazing through multi-level progressive feature disentanglement, a self-guided attention mechanism, and a disentanglement-guided module. In image dehazing tasks, self-supervised learning often need to be designed specifically for the characteristics of images under hazy conditions. To address this issue, our method leverages a generalized and effective self-supervised learning approach to enhance the model’s capability in processing hazy images.
@@ III Methodology
As shown in Fig. 2, our proposed method consists of two parts: external data augmentor and internal data augmentor. In Fig. 2 , N, C, H, and W represent the batch size, the number of channels, the height, and the width of the images, respectively. Our method utilizes the NID [13] and RSID [13], which consist of hazy images and their corresponding clear images. These datasets are used for both training and evaluating the effectiveness of the proposed method. Firstly, we propose an external- augmentation method to reduce the contrast degradation between different hazy datasets. External data augmentor makes the augmented dataset approximate the target dataset in the color space. The augmented samples obtained by our method can better match the distribution of the target domain. We believe that dehazing networks trained on a larger and more consistent set of samples are more likely to achieve superior performance.
Building upon the augmentation of different domain datasets externally, we further explore additional locally effective information by augmenting the dataset internally. The model becomes more robust by coordinating external and internal augmentation. For internal- augmentation, we introduce a weak-to-strong coordinated data augmentation self-supervised learning method to improve the local details of images. The model can focus more on the details of the images by calculating the difference between weakly and strongly augmented images.
Figure 3: A demonstration of domain gap between different dahazing datasets. We perform gamma correction on the RSID based on the RGB values of the NID. It can be observed that before correction, there is a significant difference between the distributions of the datasets in the RGB channels. After external- augmentation, the distributions of RSID images are closely aligned with those of NID images in the RGB channels.
III-A Cross-data External Alignment
We present an external-augmentor to address the domain gap between different datasets. The proposed external-augmentor achieves efficient domain adaptation without an additional training phase. Specifically, we first calculate the gamma values $\gamma_{R,G,B}$ for the three channels to be transferred to the target domain. Then, we perform gamma calibration on the images to be transferred, as follows:
@@ III-A Cross-data External Alignment
where $I_{R,G,B}$ and $O_{R,G,B}$ represent the input and output pixel intensities ([0,255]), respectively, and $\gamma$ is the gamma factor. The subscripts R, G, B represent the color channels, and the values of different channels are unique. Our method adjusts the RSID to make it more similar to the NID from color space. As shown in Fig. 3, the RSID exhibit a closer alignment with the RGB distribution of the NID after external augmentation. By performing augmentation on the target dataset, our method effectively increases the diversity of dataset and enriches it with more varied examples, which helps improve the robustness. Some of the transformed data is shown in Fig. 4. The first column shows the original RSID images, the second column shows the new RSID images after data preprocessing, and the third column shows the target dataset NID. As shown in Fig. 4, the RSID images become closer to the NID after the external augmentor. To illustrate our method intuitively, we demonstrate the algorithmic flow of external augmentation. As shown in Algorithm 1, we apply the RSID as an example to present the implementation of the channel correction. First, we calculate the average values of the three color channels (R, G, B) for both the source domain and the target domain, respectively. Then, we use the gamma correction formula to derive the required gamma value. Next, we apply the calculated gamma value to the source domain to adjust its three color channel values, making them approximate those of the target domain.
@@
Algorithm 1 External-Augmentor for Cross-data Alignment
@@
     $O_{CM-R}$=$\left(\frac{I_{R}}{255}\right)^{\frac{1}{\gamma_{R}}},O_{CM-G}$=$\left(\frac{I_{G}}{255}\right)^{\frac{1}{\gamma_{G}}},O_{CM-B}$=$\left(\frac{I_{B}}{255}\right)^{\frac{1}{\gamma_{B}}}$
@@
Algorithm 2 Internal-Augmentation with SSL
@@
III-B Feature Representation Extraction
As shown in Fig. 2, the dehazing network is composed of a 5-layer U-Net structure [11]. The network includes two downsampling and upsampling steps. In the network, the input hazy image passes through Dehaze Block modules and feature fusion modules to finally obtain the output dehazed image $\widehat{y}$. As shown in part B of Fig. 2, in the Dehaze Block module, the input feature maps undergo normalization and padding layers. Then they are projected into Q , K and V (query, key, value) using linear layers, used to calculate attention scores [12]. Additionally, the Dehaze Block module applies multi-head self-attention within a window to handle relationships between image patches. It allows the dehazing network to consider correlations between different positions while extracting features. Thus, the model can learn structural information in the image more effectively to improve the model performance.